An unknown gas at 75.1 degree C and 1.00 atm has a molar mass of 30.07 g/mol. Assuming ideal behavior, what is the density of the gas? Answer __________ g/L
Answer #1

D= masS/V

PV = nRT

n = masS/MW

PV = mass*RT/MW

then

mass/V = P*MW/(RT)

D = P*MW/(RT)

D = (1*30.07)/(0.082*(75.1+273) = 1.05345394 g/L

D = 1.05345394 g/L